Fermi National Laboratory


Accelerator Update for January 26, 2005 - January 28, 2005

Wednesday January 26
The day shift began with operations monitoring store 3945, with technicians preparing the Main Injector (MI) VCB cubical at MI-60 for removal, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

EE support technicians completed cleaning and lubricating the VCB for the quadrupole power supply Q703 at 11:11 AM.

Operations reported at 11:42 AM that the MI-60 cubical work was finished.

Operations turned on MI at 1:33 PM, but then had RF trouble, which halted stacking and MCenter beam. MI experts investigated.

At 2:46 PM, the Tevatron (TeV) suffered an abort.

Pbar began stacking at 3:14 PM.

Operations reestablished beam to MCenter at 3:24 PM.

Operators prepared the Collider Detector at Fermilab (CDF) for an access at 3:44 PM.

The Recycler began stashing antiprotons around 4 PM.

CDF came out of access at 4:30 PM.

NuMI experts completed their study at 5:30 PM.

Operations put the TeV into a dry squeeze at 7:37 PM.

Operations loaded four bunches of protons into the TeV at 8:37 PM.

Operations put the TeV into shot setup at 9:21 PM. Shot setup halts all Switchyard (SY) beam.

Operations established store 3948 at 11:59 PM with an initial luminosity of 90.04E30.

Thursday January 27
The midnight shift began with Operations monitoring store 3948, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Operations established beam to MTest at 1:40 AM.

Recycler stashed antiprotons at 6:07 AM. Pbar resumed stacking and Operations resumed sending beam to SY.

A Pbar expert began a study period at 12:33 PM.

At 2:42 PM, the MI-10 pond water pump tripped off. Technicians from the water group were doing work on variable speed pumps around the ring, but this shouldn't have caused the trip. LCW techs helped get the pump back on and stacking resumed at 3 PM.

The Pbar study ended at 7:08 PM.

Around 9:43 PM, the H- Source dome lost power. The H- source wasn't being used. Operators called an expert who suggested they temporarily power it from a local outlet, which was what they did. Repair of the source will wait until morning.

Friday January 28
The midnight shift began with Operations monitoring store 3948 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:15 AM, Recycler suffered from an emittance problem.

Operations terminated store 3948 at 4:09 AM.

Operations put the TeV into shot setup at 4:22 AM.

At 7:32 AM, Operations reported that the shot setup was on hold due to BPM problems. An expert has been helping operators resolve this problem.

The Plan for Friday and the Weekend
The plan for today is to establish a new store this morning. Pbar will then stack to 80mA and then stash antiprotons in the Recycler. There may be a TeV access on Saturday to do a small job. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/26/2005   email Fermilab